Find the volume of the solid generated by revolving the region bounded by the curves y =x2 and y =4x−x2 about the line y = 6.
1
Expert's answer
2020-06-30T17:52:53-0400

"V=\pi\int_{0}^{2}(6-y_1)^2dx-\pi\int_{0}^{2}(6-y_2)^2dx" where "y_1=x^2,y_2=4x-x^2" hence

"V=\pi\int_{0}^{2}((6-x^2)^2-(6-4x+x^2)^2)dx\approx\pi\sdot21\frac{1}{3}"